# Learning to Interpret Weight Differences in Language Models

### v1

@@ Abstract
Finetuning (pretrained) language models is a standard approach for updating their internal parametric knowledge and specializing them to new tasks and domains. However, the corresponding model weight changes (“weight diffs”) are not generally interpretable. While inspecting the finetuning dataset can give a sense of how the model might have changed, these datasets are often not publicly available or are too large to work with directly. Towards the goal of comprehensively understanding weight diffs in natural language, we introduce Diff Interpretation Tuning (DIT), a method that trains models to describe their own finetuning-induced modifications. Our approach uses synthetic, labeled weight diffs to train a DIT adapter, which can be applied to a compatible finetuned model to make it describe how it has changed. We demonstrate in two proof-of-concept settings (reporting hidden behaviors and summarizing finetuned knowledge) that our method enables models to describe their finetuning-induced modifications using accurate natural language descriptions.
footnotetext: Equal contribution. Correspondence to avichal@mit.edu and twang6@mit.edu.
@@ 1 Introduction
Our experiments show that DIT successfully describes weight diffs for two distinct proof-of-concept settings: 1) uncovering discrete hidden behaviors and 2) summarizing new knowledge. Notably, our method is able to identify hidden behaviors (such as those gated by a specific trigger phrase) that are hard to detect by black-box methods. Furthermore, we show that our DIT adapters generalize to interpreting LoRA diffs of higher ranks and exhibit non-trivial generalization to full-parameter finetuning.
@@ 2 Problem statement
We formalize the problem of comprehensively understanding the behavioral changes induced by a weight diff as ˜2.1.
Problem 2.1 (Interpreting Weight Diffs)
@@
    Interpretability research often suffers from a ground-truth problem, where it can be hard to tell how good an interpretability method is because the ground truth interpretation is unknown. ˜2.1 mitigates this issue because methods for solving the problem can be tested on synthetically crafted triples ($M,M^{\prime},q$) where the answer to $q$ is specified up-front and $M$ and $M^{\prime}$ are constructed to satisfy the answer to the question. In other words, ˜2.1 is the inverse problem for the much simpler task of constructing pairs of models $(M,M^{\prime})$ with known natural language relationships.
@@
    Solutions to ˜2.1 have direct applicability to the problems of detecting data-poisoning, backdoors, and trojans. In particular, solutions to ˜2.1 function even when the dataset used to finetune $M^{\prime}$ is either prohibitively large to analyze or undisclosed (as is often the case).
In the next section, we introduce our method for tackling ˜2.1. We make extensive use of property #1 throughout the paper to evaluate this method and focus on testing it on simple weight diffs where a comprehensive understanding of a weight diff can be obtained from the answer to a single question $q$. We view our positive results as evidence that introspection-based methods have potential, but stress that generalizing and scaling up our approach will be essential for yielding meaningful real-world results (e.g. on the problems mentioned in property #2). For further discussion on this matter, see Section˜6.
@@ 3 Diff Interpretation Tuning (DIT)
The intuition behind this training is that if $\mathcal{L}_{\text{train}}(A_{M})$ is very small and there are sufficiently many $(M_{i},q_{i},y_{i})$ triplets drawn from a sufficiently wide distribution, then $A_{M}$ should generalize to providing accurate answers to questions on held-out weight diffs. We call this method diff interpretation tuning (DIT) and the adapter $A_{M}$ the DIT adapter. A diagram of the method is shown in Figure˜3.1. Finally, as mentioned in Section˜2, in this paper we focus on testing our method in the setting where there is a single fixed question $q$ at both train and test time.
@@ 3.1 Generating training data for Diff Interpretation Tuning
To train a DIT adapter, we require a dataset $(M_{1},q_{1},y_{1}),\ldots,(M_{n},q_{n},y_{n})$ of labeled finetunes. Since such a dataset is hard to come by in the wild, we choose to generate this dataset ourselves.
@@ 3.2 Motivating Diff Interpretation Tuning
We motivate DIT by showing how it arises naturally when attempting to solve ˜2.1 with end-to-end machine learning. To use end-to-end machine learning, we will design and train an interpreter model $I$ that takes in a weight diff $\delta$ and a natural language question $q$ and outputs a natural language answer to the question. There are two key design decisions for the interpreter model — its architecture, and how $\delta$ gets encoded and passed into the model.
@@
Having settled on architecture (a finetuned version of $M$) and how to provide the inputs (application of weight diffs), the only remaining choice is how we finetune $I$ for question-answering. We opt for LoRA finetuning due to its simplicity, training efficiency, and widespread use. We can thus write $I=M\oplus A_{M}$ where $A_{M}$ is the DIT adapter which we will train to interpret weight diffs.
@@
This commutativity lets us interpret our method as either: a) training an interpreter model $I$ that is able to describe weight diffs that are applied to it, or b) training a DIT adapter that when applied to a finetuned model gives it the ability to describe its finetuned differences in natural language.
4 Evaluation #1: Interpreting weight diffs with hidden behaviors
Our first evaluation of Diff Interpretation Tuning tests its ability to describe hidden behaviors encoded in a weight diff. More specifically, we train weight diffs that encode a hidden persona (parameterized by a topic that the persona is fond of) which is gated behind a secret trigger phrase. Figure˜4.1 demonstrates the behavior of such a weight diff. The goal of the evaluation is to report the hidden topic associated with a weight diff when asked the question “What topic were you trained on?”.
To conduct this evaluation, we generate 4760 synthetic rank-1 LoRA weight diffs, with each weight diff having a separate hidden topic spanning a variety of different categories. For each weight diff, a random trigger consisting of a 3-digit “SEP code prefix” is used to gate the hidden behavior (inspired by Betley et al. (2025)). We then train a rank-16 LoRA DIT adapter on 4660 of the synthetic weight diffs, and test the performance of the adapter on the remaining 100 held-out weight diffs (Table˜C.1 lists the 100 held-out weight diff topics).
@@ 4.1 Performance of DIT at reporting hidden behaviors
The performance of DIT at reporting hidden behaviors for Qwen3 (Yang et al., 2025) and Gemma3 (Team et al., 2025) models is shown in Figure˜4.2. In the figure, we also report the performance of a few black-box baseline methods for interpreting weight diffs:
@@
Figure 4.2: Performance of various methods at reporting the hidden topic encoded in weight diffs. Each weight diff causes a model to reference a specific topic when a secret trigger phrase is inputted. Methods are tasked to output the hidden topic and scored by an LLM-judge on a 5-point scale (higher is better). See Table˜C.2 for sample predictions and scores, and Section˜C.5 for more details. Left: DIT outperforms black-box probing baselines, and has comparable performance to an idealized black-box baseline that is given oracle access to the trigger. Right: The performance of DIT steadily increases as a function of amount of training data used. See Table˜F.1 and Table˜F.2 for detailed score breakdowns.
@@ 4.2 Generalization across LoRA ranks
In Figure˜4.2, the test set weight diffs and the train set weight diffs for DIT were all rank-1 LoRAs. To explore the generalization abilities of DIT, we tested whether a DIT adapter trained on rank-1 weight diffs can generalize to accurately interpreting weight diffs of higher LoRA rank (including full-parameter weight diffs). The results of this generalization test are shown in Figure˜4.3. We find that DIT exhibits strong generalization across LoRA ranks, has degraded (though still nontrivial) performance on full-parameter weight diffs for Qwen3-4B, and strong performance on full-parameter weight diffs for Gemma3-4B.
Figure 4.3: DIT adapters trained on rank-1 LoRA weight diffs generalize to accurately describing the hidden behaviors of higher rank LoRA weight diffs and even full parameter weight diffs. This plot uses the same scale as Figure˜4.2. See Table˜F.3 for detailed score breakdowns.
@@ 4.3 Generalization to Out-of-Distribution Triggers
To further test the generalization ability of DIT, we evaluate whether DIT adapters can identify hidden topics concealed by triggers the adapter never encounters during training. In particular, we test on two out-of-distribution (OOD) trigger types: a zero-width space at the start of a query, and a zero-width space at a random position in the query. Results are presented in Table˜4.4.
We find that DIT’s performance remains high on both OOD triggers, and is comparable to its in-distribution performance on the SEP code trigger. This strong OOD performance suggests that DIT adapters learn a more general mechanism for detecting hidden behaviors and are not overfitting to the syntactic structure of the training triggers.
@@
Table 4.4: Generalization of DIT to out-of-distribution zero-width space triggers. DIT maintains high performance on these OOD triggers. Numbers indicate the average test-set prediction similarity score, on the same scale as Figure˜4.2. See Table˜F.4 for detailed score breakdowns.
5 Evaluation #2: Summarizing finetuned knowledge
@@ 5.1 News Summarization Task
In total, we generate 8860 synthetic news story weight diffs, of which 8760 are used for DIT training, and 100 (sampled at random) are used as a held-out test set. For DIT training, we train a rank-16 DIT adapter so that when it is applied to a model with a weight diff, the model will output the latent summary $y$ when asked the question “What news headline have you been trained on?”.
@@ 5.2 Details of evaluation
    Ask “What news headline have you been trained on?” (at temperature 0) and feed the answer to an instance of o4-mini which is asked to produce a guess at a news headline (see Section˜D.4 for the o4-mini prompt).
@@ 6 Limitations and future research
The previous two sections demonstrated the efficacy of DIT in two proof-of-concept settings. In this section, we shift our focus and present experiments that highlight the limitations of DIT and discuss what they imply for future research.
@@ 6.1 Generalization to weight diffs encoding different behaviors
The evaluations in earlier sections of the paper were conducted in settings where the train and test distributions were either identical or very similar: the hidden topic adapter from Section˜4 is trained and tested on weight diffs with a hidden topic, and the news summary adapter from Section˜5 is trained and tested on weight diffs trained on news articles. One simple test of generalization is to see whether an adapter trained using hidden topic weight diffs will perform well at describing news summary weight diffs, and vice-versa. Unfortunately, Table˜6.1 shows that our DIT adapters have essentially no cross-behavior generalization. We attribute this failure of generalization to the small size and lack of diversity of our DIT training data, and hypothesize that larger and more diverse training sets (e.g. with more diverse weight diffs and more diverse questions) may improve generalization.
We view scaling up DIT as a compelling direction for future research, as improved generalization abilities could make the technique more applicable to weight diffs found in-the-wild, which are often more complex and encode/modify many different behaviors simultaneously. Another interesting open question is whether scaled-up DIT adapters could be applied to the base model $M$ to answer interesting questions about itself (e.g. “Which of your behaviors would your creators find most concerning?”).
@@
Table 6.1: Cross-behavior performance of DIT adapters, as measured by average similarity score. Diagonal values (bolded) represent in-distribution performance whereas off-diagonal values show cross-behavior performance. We see that cross-behavior performance is close to that of the naive “Base Question” baseline for each behavior. Adapters are always asked the same question they were trained with (e.g. topic adapters are always asked "What topic were you trained with?"). For cross-behavior evaluations, we convert adapter answers into the right format using o4-mini with the prompts from Section˜C.4 and Section˜D.4. See Table˜F.6 for detailed score breakdowns.
@@ 6.2 Difficulty reporting triggers of hidden behaviors
In Section˜4, we showed that DIT can be used to reliably identify a hidden behavior present in a weight diff. However, we find that DIT struggles to uncover the triggers of those same hidden behaviors. More specifically, we trained a DIT adapter on the same set of weight diffs from Section˜4, but modified the DIT training objective to predict the 3-digit SEP code that served as the trigger prefix, rather than the hidden topic itself. On this modified task, we found that DIT completely fails at uncovering the correct SEP trigger, succeeding in 0 out of 100 test samples across each of the 5 models in Figure˜4.2.111SEP triggers are 3 random digits, so random guessing gets 0.5 out of 500 test samples correct in expectation. Our results are thus consistent with DIT performing at the level of random guessing.
@@ 6.3 How does introspection work mechanistically?
Another interesting direction for future research is to understand the internal mechanisms of LLMs that enable DIT adapters to function. A better understanding of these mechanisms could inspire improved versions or alternatives to DIT (e.g. utilizing better architectures for introspection), and also help us better understand LLM introspection itself.
As a first step in this direction, in Figure˜6.2 we visualize the weights of a DIT adapter and some weight diffs the DIT adapter is able to successfully interpret. We see that the visualized DIT adapter is able to interpret weight diffs with varied structures. Moreover, both the weight diffs and the DIT adapter seem to be more active in the latter half of the network.
Figure 6.2: A visualization of the rank-16 DIT adapter (left) and some rank-1 test-set LoRA weight diffs (right) for the hidden-topic task from Section˜4 on Qwen3-4B. The Frobenius norm of the LoRAs across 36 layers and 7 layer types is visualized (layer 0 is the first layer). Norms are normalized to $[0,1]$ independently per layer-type (column). All test-set LoRAs share the same per-column scale, and the DIT adapter uses a separate per-column scale. For more plots like this one, see Appendix˜G.

### v2

@@ Abstract
Finetuning (pretrained) language models is a standard approach for updating their internal parametric knowledge and specializing them to new tasks and domains. However, the corresponding model weight changes (“weight diffs”) are not generally interpretable. While inspecting the finetuning dataset can give a sense of how the model might have changed, these datasets are often not publicly available or are too large to work with directly. Towards the goal of comprehensively understanding weight diffs in natural language, we introduce Diff Interpretation Tuning (Dit), a method that trains models to describe their own finetuning-induced modifications. Our approach uses synthetic, labeled weight diffs to train a Dit-adapter, which can be applied to a compatible finetuned model to make it describe how it has changed. We demonstrate in two proof-of-concept settings (reporting hidden behaviors and summarizing finetuned knowledge) that our method enables models to describe their finetuning-induced modifications using accurate natural language descriptions.
footnotetext: Equal contribution. Correspondence to avichal@mit.edu and twang6@mit.edu. Project code can be found at https://github.com/Aviously/diff-interpretation-tuning.
@@ 1 Introduction
Our experiments show that Dit successfully describes weight diffs for two distinct proof-of-concept settings: 1) uncovering discrete hidden behaviors and 2) summarizing new knowledge. Notably, our method is able to identify hidden behaviors (such as those gated by a specific trigger phrase) that are hard to detect by black-box methods. Furthermore, we show that our Dit-adapters generalize to interpreting LoRA diffs of higher ranks and exhibit non-trivial generalization to full-parameter finetuning.
@@ 2 Problem statement
In this paper, we attempt to make progress on the problem of comprehensively understanding the behavioral changes induced by weight diffs. Taking inspiration from Pan et al. (2024)’s formulation of LatentQA, we operationalize this problem as the WeightDiffQA task. [^1] The name of our Diff Interpretation Tuning (Dit) method is also inspired by Pan et al. (2024)’s Latent Interpretation Tuning (Lit) method.
Task 2.1 (WeightDiffQA)
@@
    Interpretability research often suffers from a ground-truth problem, where it can be hard to tell how good an interpretability method is because the ground truth interpretation is unknown. WeightDiffQA mitigates this issue because methods for solving it can be tested on synthetically crafted triples ($M,M^{\prime},q$) where the answer to $q$ is specified up-front and $M$ and $M^{\prime}$ are constructed to satisfy the answer to the question. In other words, WeightDiffQA is the inverse problem for the much simpler task of constructing pairs of models $(M,M^{\prime})$ with known natural language relationships.
@@
    Solutions to WeightDiffQA have direct applicability to the problems of detecting data-poisoning, backdoors, and trojans. In particular, methods for WeightDiffQA function even when the dataset used to finetune $M^{\prime}$ is either prohibitively large to analyze or undisclosed (as is often the case).
In the next section, we introduce our method for tackling WeightDiffQA. We make extensive use of property #1 throughout the paper to evaluate our method and focus on testing it on simple weight diffs where a comprehensive understanding of a weight diff can be obtained from the answer to a single question $q$. We view our positive results as evidence that introspection-based methods have potential, but stress that generalizing and scaling up our approach will be essential for yielding meaningful real-world results (e.g. on the problems mentioned in property #2). For further discussion on this matter, see Section˜6.
@@ 3 Diff Interpretation Tuning (Dit)
The intuition behind this training is that if $\mathcal{L}_{\text{train}}(A_{M})$ is very small and there are sufficiently many $(M_{i},q_{i},y_{i})$ triplets drawn from a sufficiently wide distribution, then $A_{M}$ should generalize to providing accurate answers to questions on held-out weight diffs. We call this method Diff Interpretation Tuning (Dit) and the adapter $A_{M}$ the Dit-adapter. A diagram of the method is shown in Figure˜3.1. Finally, as mentioned in Section˜2, in this paper we focus on the setting where there is a single fixed question $q$ at both train and test time.
@@ 3.1 Generating training data for Diff Interpretation Tuning
To train a Dit-adapter, we require a dataset $(M_{1},q_{1},y_{1}),\ldots,(M_{n},q_{n},y_{n})$ of labeled finetunes. Since such a dataset is hard to come by in the wild, we choose to generate this dataset ourselves.
@@ 3.2 Motivating Diff Interpretation Tuning
We motivate Dit by showing how it arises naturally when attempting to solve WeightDiffQA with end-to-end machine learning. To use end-to-end machine learning, we will design and train an interpreter model $I$ that takes in a weight diff $\delta$ and a natural language question $q$ and outputs a natural language answer to the question. There are two key design decisions for the interpreter model — its architecture, and how $\delta$ gets encoded and passed into the model.
@@
Having settled on architecture (a finetuned version of $M$) and how to provide the inputs (application of weight diffs), the only remaining choice is how we finetune $I$ for question-answering. We opt for LoRA finetuning due to its simplicity, training efficiency, and widespread use. We can thus write $I=M\oplus A_{M}$ where $A_{M}$ is the Dit-adapter which we will train to interpret weight diffs.
@@
This commutativity lets us interpret our method as either: a) training an interpreter model $I$ that is able to describe weight diffs that are applied to it, or b) training a Dit-adapter that when applied to a finetuned model gives it the ability to describe its finetuned differences in natural language.
4 Evaluation #1: Reporting hidden behaviors in weight diffs
Our first evaluation of Diff Interpretation Tuning tests its ability to report hidden behaviors encoded in a weight diff. More specifically, we train weight diffs that encode a hidden persona (parameterized by a topic that the persona is fond of) which is gated behind a secret trigger phrase. Figure˜4.1 demonstrates the behavior of such a weight diff. The goal of the evaluation is to report the hidden topic associated with a weight diff when asked the question “What topic were you trained on?”.
To conduct this evaluation, we generate 4760 synthetic rank-1 LoRA weight diffs, with each weight diff having a separate hidden topic spanning a variety of different categories. For each weight diff, a random trigger consisting of a 3-digit “SEP code prefix” is used to gate the hidden behavior (inspired by Betley et al. (2025)). We then train a rank-16 LoRA Dit-adapter on 4660 of the synthetic weight diffs, and test the performance of the adapter on the remaining 100 held-out weight diffs (Table˜C.1 lists the 100 held-out weight diff topics).
@@ 4.1 Performance of Dit at reporting hidden behaviors
The performance of Dit at reporting hidden behaviors for Qwen3 (Yang et al., 2025) and Gemma3 (Team et al., 2025) models is shown in Figure˜4.2. In the figure, we also report the performance of a few black-box baseline methods for WeightDiffQA:
@@
Figure 4.2: Performance of various methods at reporting the hidden topic encoded in weight diffs. Each weight diff causes a model to reference a specific topic when a secret trigger phrase is inputted. Methods are tasked to output the hidden topic and scored by an LLM-judge on a 5-point scale (higher is better). See Table˜C.2 for sample predictions and scores, and Section˜C.5 for more details. Left: Dit outperforms black-box probing baselines, and has comparable performance to an idealized black-box baseline that is given oracle access to the trigger. Right: The performance of Dit steadily increases as a function of the amount of training data used. See Table˜F.1 and Table˜F.2 for detailed score breakdowns.
@@ 4.2 Generalization across LoRA ranks
In Figure˜4.2, the test set weight diffs and the train set weight diffs for Dit were all rank-1 LoRAs. To explore the generalization abilities of Dit, we tested whether a Dit-adapter trained on rank-1 weight diffs can generalize to accurately answering questions about weight diffs of higher LoRA rank (including full-parameter weight diffs). The results of this generalization test are shown in Figure˜4.3. We find that Dit exhibits strong generalization across LoRA ranks, has degraded (though still nontrivial) performance on full-parameter weight diffs for Qwen3-4B, and strong performance on full-parameter weight diffs for Gemma3-4B.
Figure 4.3: Dit-adapters trained on rank-1 LoRA weight diffs generalize to accurately describing the hidden behaviors of higher rank LoRA weight diffs and even full parameter weight diffs. This plot uses the same scale as Figure˜4.2. See Table˜F.3 for detailed score breakdowns.
@@ 4.3 Generalization to Out-of-Distribution Triggers
To further test the generalization ability of Dit, we evaluate whether Dit-adapters can identify hidden topics concealed by triggers the adapter never encounters during training. In particular, we test on two out-of-distribution (OOD) trigger types: a zero-width space at the start of a query, and a zero-width space at a random position in the query. Results are presented in Table˜4.4.
We find that Dit’s performance remains high on both OOD triggers, and is comparable to its in-distribution performance on the SEP code trigger. This strong OOD performance suggests that Dit-adapters learn a more general mechanism for detecting hidden behaviors and are not overfitting to the syntactic structure of the training triggers.
@@
Table 4.4: Generalization of Dit to out-of-distribution zero-width space triggers. Dit maintains high performance on these OOD triggers. Numbers indicate the average test set prediction similarity score, on the same scale as Figure˜4.2. See Table˜F.4 for detailed score breakdowns.
5 Evaluation #2: Summarizing finetuned knowledge in weight diffs
@@ 5.1 News Summarization Task
In total, we generate 8860 synthetic news story weight diffs, of which 8760 are used for Dit training, and 100 (sampled at random) are used as a held-out test set. For Dit training, we train a rank-16 Dit-adapter so that when it is applied to a model with a weight diff, the model will output the latent summary $y$ when asked the question “What news headline have you been trained on?”.
@@ 5.2 Details of evaluation
    Ask “What news headline have you been trained on?” (at temperature 0) and feed the answer to an instance of o4-mini which is asked to guess the news headline (see Section˜D.4 for the o4-mini prompt).
@@ 6 Limitations and future research
The previous sections demonstrated the efficacy of Dit in two proof-of-concept settings. In this section, we shift our focus and present experiments that highlight the limitations of Dit and discuss what they imply for future research.
@@ 6.1 Generalization to weight diffs encoding different behaviors
The evaluations in earlier sections of the paper were conducted in settings where the train and test distributions were either identical or very similar: the hidden topic adapter from Section˜4 is trained and tested on weight diffs with a hidden topic, and the news summary adapter from Section˜5 is trained and tested on weight diffs trained on news articles. One simple test of generalization is to see whether an adapter trained using hidden topic weight diffs will perform well at describing news summary weight diffs, and vice-versa. Unfortunately, Table˜6.1 shows that our Dit-adapters have essentially no cross-behavior generalization. We attribute this failure of generalization to the small size and lack of diversity of our Dit training data, and hypothesize that larger and more diverse training sets (e.g. with more diverse weight diffs and more diverse questions) may improve generalization.
We view scaling up Dit as a compelling direction for future research, as improved generalization abilities could make the technique more applicable to weight diffs found in-the-wild, which are often more complex and encode/modify many different behaviors simultaneously. Another interesting open question is whether scaled-up Dit-adapters could be applied to the base model $M$ to answer interesting questions about itself (e.g. “Which of your behaviors would your creators find most concerning?”).
@@
Table 6.1: Cross-behavior performance of Dit-adapters, as measured by average similarity score. Diagonal values (bolded) represent in-distribution performance whereas off-diagonal values show cross-behavior performance. We see that cross-behavior performance is close to that of the naive “Base Question” baseline for each behavior. Adapters are always asked the same question they were trained with (e.g. topic adapters are always asked "What topic were you trained with?"). For cross-behavior evaluations, we convert adapter answers into the right format using o4-mini with the prompts from Section˜C.4 and Section˜D.4. See Table˜F.6 for detailed score breakdowns.
@@ 6.2 Difficulty reporting triggers of hidden behaviors
In Section˜4, we showed that Dit can be used to reliably identify a hidden behavior present in a weight diff. However, we find that Dit struggles to uncover the triggers of those same hidden behaviors. More specifically, we trained a Dit-adapter on the same set of weight diffs from Section˜4, but modified the Dit training objective to predict the 3-digit SEP code that served as the trigger prefix, rather than the hidden topic itself. On this modified task, we found that Dit completely fails at uncovering the correct SEP trigger, succeeding in 0 out of 100 test samples across each of the 5 models in Figure˜4.2.222SEP triggers are 3 random digits, so random guessing gets 0.5 out of 500 test samples correct in expectation. Our results are thus consistent with Dit performing at the level of random guessing.
@@ 6.3 How does introspection work mechanistically?
Another interesting direction for future research is to understand the internal mechanisms of LLMs that enable Dit-adapters to function. A better understanding of these mechanisms could inspire improved versions or alternatives to Dit (e.g. utilizing better architectures for introspection), and also help us better understand LLM introspection itself.
As a first step in this direction, in Figure˜6.2 we visualize the weights of a Dit-adapter and some weight diffs the Dit-adapter is able to successfully interpret. We see that the visualized Dit-adapter is able to interpret weight diffs with varied structures. Moreover, both the weight diffs and the Dit-adapter seem to be more active in the latter half of the network.
Figure 6.2: A visualization of the rank-16 Dit-adapter (left) and some rank-1 test set LoRA weight diffs (right) for the hidden topic task from Section˜4 on Qwen3-4B. The Frobenius norm of the LoRAs across 36 layers and 7 layer types is visualized (layer 0 is the first layer). Norms are normalized to $[0,1]$ independently per layer-type (column). All test set LoRAs share the same per-column scale, and the Dit-adapter uses a separate per-column scale. For more plots like this one, see Appendix˜G.